What is the implied interest rate for the first choice? _______% and which investment should you choose? A or B___________________ Why? __________________________



A. You can invest $1000 today and receive $1158 in 3 years. The investment is considered low risk.                


B. You can invest the $1000 in a bank account paying 5% for 5 years.     



What is the implied interest rate for the first choice? _______%   and which investment should you choose?  A or B___________________   Why? __________________________
Answer #1

What is the implied interest rate for the first choice?

= (FV/ PV)1/n - 1 = (1,158 / 1,000)1/3 - 1 =  5.01%

and which investment should you choose?

A

Why? Because

  • All other factors being nearly constant, in A, the tenure is only 3 years. You can reinvest the proceeds in any better opportunity then. If there is no better opportunity, you can always roll it over for one more term of 3 years.